A wire with length 15m and diameter 1mm has resistance 0.28Ω. What is the material this wire is made of?


1
Expert's answer
2022-01-10T09:17:08-0500

Given:

"l=15\rm\: m"

"d=1\rm\: mm"

"R=0.28\rm\: \Omega"


The resistance of a wire

"R=\rho\frac{l}{A}=\rho\frac{4l}{\pi d^2}"

Hence, the resistivity of a material

"\rho=R\frac{\pi d^2}{4l}\\
=0.28*\frac{3.14*0.001^2}{4*15}=1.5*10^{-8}\:\rm \Omega\cdot m"

The material is silver.